120 grams of sodium hydroxide completely reacts with phosphoric acid (H3PO4) in the reaction 3 NaOH + H3PO4 -> Na3PO4 + 3 H2O what mass of water is produced?


1
Expert's answer
2021-06-03T05:28:05-0400

Moles of Sodium hydroxide

= 120 / 40 =3 mol.


3 mol. Of NaOH gives 3 mol . Of water


Mass of water formed = 3 × 18 = 54 g .